delaneyglover765:

Find the percent of markup. Round to the nearest tenth. $35.50 to $70.00

Bigbosssaint21:

Are there answer options?

Tranquility:

Hello! Welcome to QuestionCove! To calculate the percent of markup, use the following formula Markup % = (Price - Cost)/Price * 100 For example, if we are selling something for $100 but it costs $50, then the percent markup is Markup % = (100-50)/100 * 100 = 50/100 * 100 = 50% markup

Tranquility:

In your question, $35.50 is the cost and we're raising the price to $70.00 Should be an easy plug and chug